OpenStudy (jotopia34):

How do you take the derivative of R(w)=3^wlog(w)? Also can you tell me how you can have just log(w) and not a log with a base?

OpenStudy (anonymous):

\[\log_{10} w=\log_{}w \] base is 10 \[R^'(x)=3^x \log(x) \log (x) +\frac{ 1 }{ x }3^x\]

OpenStudy (jotopia34):

so if there is no base written, 10 is assumed?

OpenStudy (anonymous):

yes

OpenStudy (jotopia34):

thx!!

OpenStudy (amistre64):

if there is no written base, then you cant really make an assumption of 10 unless the text or material that you are using defines it as such. http://www.wolframalpha.com/input/?i=logarithm&lk=1&a=ClashPrefs_*MathWorld.Logarithm-